# Radiative Corrections to Light Thermal Pseudo-Dirac Dark Matter

### v1

@@ III.3 Relic Density
Figure 4: Value of $\varepsilon$ producing the observed relic abundance of dark matter as a function of $m_{\chi}$, for $\alpha_{D}=$ 0.05, 0.3 and 0.5 (red, blue and magenta lines respectively) based on LO only (solid lines) and NLO + LO (dashed lines) cross sections and for the indicated value of $m_{A^{\prime}}$ on each panel. The dark Higgs mass is chosen to be $\geq m_{A^{\prime}}$, such that its contribution is negligible. For the $m_{A^{\prime}}=3m_{\chi}$ panel, the gray shaded regions are the exclusions from past accelerator experiments, and the green, purple and orange dotted lines are projections from the upcoming LDMX, $M_{3}$ and Belle II experiments respectively Gori et al. (2022).
In Fig. 4 we show plots of kinetic mixing as a function of DM mass for a few representative values of the ratio of $m_{A^{\prime}}/m_{\chi}$, obtained by scanning through the kinetic mixing for each DM mass, finding the combination resulting in the observed relic abundance $\Omega h^{2}=0.12$ Aghanim et al. (2018). Each plot shows the required kinetic mixing parameter for three different values of the dark sector coupling $\alpha_{D}$, corresponding to the LO only (solid lines) and NLO + LO (dashed lines) computations respectively. For $m_{A^{\prime}}/m_{\chi}=3$, which is typically chosen as a benchmark parameter point in the literature, we also indicate the current experimental constraints (a combination of both accelerator and astrophysical probes Gori et al. (2022)) as the gray shaded region. In that panel, green, purple and orange dotted lines are projections from the upcoming LDMX, $M_{3}$ and Belle II experiments, respectively.
@@ IV Dark matter direct detection
Following ref. Hill and Solon (2015), we factorize the matrix element as:
$|\overline{\mathcal{M}_{EFT}}|^{2}$ $=$ $|c_{EFT}|^{2}\frac{1}{4}\sum_{spin}(J_{\chi}^{\mu}J^{f}_{\mu})^{\dagger}(J_{\chi}^{\mu}J^{f}_{\mu}),$ (13)
where $J_{\chi}^{\mu}=\bar{u}_{\chi}\gamma^{\mu}u_{\chi}$ and $J^{f}_{\mu}=\bar{u}_{f}\gamma_{\mu}u_{f}$ represent the dark sector and SM currents respectively. Here $f$ can represent either electrons or nucleons. The Wilson coefficient $c_{EFT}$ includes contributions from the tree level dark photon exchange, plus NLO corrections from the vertex correction, DM self-energy and the dark photon vacuum polarization, analogous to those represented in Fig. 2. Given our focus on sub-GeV DM and the prospects for its detection at upcoming experiments, we restrict our attention to DM scattering with electrons. However our general computations can be extended to nucleon scattering with the incorporation of the appropriate form factors for either spin-dependent or spin-independent scattering.
The differential DM-electron scattering cross-section as a function of the momentum transfer is usually written
@@
$\overline{\sigma_{e}}=\overline{\sigma}_{e}^{\rm LO}+\overline{\sigma}_{e}^{\rm NLO}+{\cal O}\left(\alpha_{D}^{3}\right).$ (15)
@@
$\overline{\sigma}_{e}^{\rm NLO}$ $=\frac{\alpha_{D}\,\overline{\sigma}_{e}^{\rm LO}}{6\,\pi\,m_{A^{\prime}}^{2}\,m_{\chi}^{2}}\Bigg{\{}m_{A^{\prime}}^{2}\left(m_{A^{\prime}}^{2}+2m_{\chi}^{2}\right)\Bigg{(}3B_{0}\left(m_{\chi}^{2};m_{A^{\prime}}^{2},m_{\chi}^{2}\right)-3\left(m_{A^{\prime}}^{2}-2m_{\chi}^{2}\right)C_{0}\left(m_{\chi}^{2},m_{\chi}^{2},0;m_{\chi}^{2},m_{A^{\prime}}^{2},m_{\chi}^{2}\right)$ (17)
$+2m_{\chi}^{2}\left(2B^{\prime}_{0}\left(m_{A^{\prime}}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)+3B^{\prime}_{0}\left(m_{\chi}^{2};m_{A^{\prime}}^{2},m_{\chi}^{2}\right)\right)\Bigg{)}-8m_{\chi}^{4}B_{0}\left(m_{A^{\prime}}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)$
$+\left(-3m_{A^{\prime}}^{4}-6m_{A^{\prime}}^{2}m_{\chi}^{2}+8m_{\chi}^{4}\right)B_{0}\left(0;m_{\chi}^{2},m_{\chi}^{2}\right)\Bigg{\}}$
@@
Figure 5: Dark matter direct detection parameter space. The left panel shows the ratio of NLO to LO electron scattering cross-section vs $m_{A^{\prime}}/m_{\chi}$ ratio for $m_{\chi}=1$ MeV (red, dashed line) and $m_{\chi}=1$ GeV (green solid line). The right panel shows dark matter-electron scattering cross-section vs $m_{\chi}$ for dark matter thermal relic abundance with $\alpha_{D}=$, 0.05, 0.3 and 0.5, represented by the red, blue and magenta lines respectively. We show both the LO only (solid lines) and NLO + LO (dashed lines) for the $m_{A^{\prime}}/m_{\chi}=2$ case. The upper gray shaded region represents the current model independent direct detection constraints from a combination of the XENON1T, PANDAX and SENSEI experiments and the vertical gray shaded region is the constraint from $\Delta N_{eff}$. All constraints were obtained from ref. Essig et al. (2022).
@@
In Fig. 5, we show the direct detection constraints on the model. The left panel characterizes the importance of the one-loop correction, parameterized by the quantity $\overline{\sigma}_{eNLO}/(\alpha_{D}\overline{\sigma}_{eLO})$ in terms of the ratio $m_{A^{\prime}}/m_{\chi}$ for 1 GeV and 1 MeV DM masses.
@@
In the right panel of the figure, we show the averaged DM-electron scattering cross-section vs DM mass. The red, blue and magenta lines represent DM produced through the thermal freeze-out mechanism for benchmark choices of $\alpha_{D}=$ 0.05, 0.3 and 0.5 respectively. The dashed lines are NLO + LO contributions, while the solid lines are only at LO. For illustration, we only show the case for $m_{A^{\prime}}/m_{\chi}\simeq$ 2. We see a clear dependence on the dark sector coupling $\alpha_{D}$ and differences between the NLO + LO and LO only contributions.
@@
Also shown as the grey shaded region are the model independent constraints on this parameter space. Below $\sim$ 10 MeV, there is a stringent bound from big bang nucleosynthesis and $\Delta N_{eff}$, whereas the upper grey shaded region indicates bounds from a combination of direct detection experiments, including XENON1T, PANDAX and SENSEI (obtained from ref. Essig et al. (2022)).
@@ V Discussion and Conclusions
In this work, we revisit the computations for sub-GeV thermal dark matter processes. We consider a highly motivated and sought after scenario, in which (pseudo)-Dirac dark matter couples to the SM via a massive dark photon, representing a compelling benchmark model of dark matter. We focus on the corrections at order $\alpha_{D}$, which is typically invoked to be large in order to obtain the observed relic density of the dark matter.
We compute the NLO corrections for thermal (pseudo)-Dirac dark matter annihilating in the early universe, as this represents a very important target for upcoming experiments. We find that the NLO corrections can be as large as 10’s of percents, depending on the $m_{A^{\prime}}/m_{\chi}$ ratio and the strength of the dark sector coupling constant $\alpha_{D}$. While this is not a dramatic difference compared to the targets inferred from tree level calculations, it is an important result to establish the ability of proposed experiments to discover or constrain MeV scale dark matter, and to precisely quantify what we will learn from such experiments in the future. Furthermore, our results show that factorization into the “yield parameter” $Y$, as is common practise in the literature, can be misleading, especially in the resonance region, where the NLO corrections are more significant.
For completeness, we also show the direct detection limits based on DM-electron scattering. We find that the NLO corrections are slightly more visible in the resonance region and become less significant for larger mass ratios (similar to the annihilation case). We leave a more dedicated study of pseudo-Dirac dark matter with larger mass splittings as well as other Lorentz structures of interest to future studies.
@@ Acknowledgements
We thank Basudeb Dasgupta for useful discussions. GM acknowledges support from the UC office of the President through the UCI Chancellor’s Advanced Postdoctoral Fellowship. GM is also supported in part by the National Sciences and Engineering Research Council of Canada. TMPT and GM are supported in part by the U.S. National Science Foundation under Grant PHY-2210283. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
@@ B.2 Dark Higgs Contribution
$\Pi_{1}(k^{2})$ $=\frac{\alpha_{D}}{16\pi(\epsilon-3)k^{2}}\Bigg{[}\Bigg{(}-2k^{2}\left(m_{A^{\prime}}^{2}(2\,\epsilon-5)+m_{H}^{2}\right)$ (28)
$+k^{4}+\left(m_{A^{\prime}}^{2}-m_{H}^{2}\right)^{2}\Bigg{)}B_{0}\left(k^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)$
$+A_{0}\left(m_{H}^{2}\right)\left((23-8\,\epsilon)k^{2}+m_{A^{\prime}}^{2}-m_{H}^{2}\right)$
$-\left(A_{0}\left(m_{A^{\prime}}^{2}\right)\left(k^{2}+m_{A^{\prime}}^{2}-m_{H}^{2}\right)\right)\Bigg{]}$
@@
$\delta m_{A^{\prime}}^{2}$ $=\frac{\alpha_{D}}{16\pi m_{A^{\prime}}^{2}(\epsilon-3)}\,\text{Re}\Bigg{[}\Bigg{(}-4m_{A^{\prime}}^{4}(\epsilon-3)$ (29)
@@
$-A_{0}\left(m_{H}^{2}\right)\left(8m_{A^{\prime}}^{2}(\epsilon-3)+m_{H}^{2}\right)$
$+\left(m_{H}^{2}-2m_{A^{\prime}}^{2}\right)A_{0}\left(m_{A^{\prime}}^{2}\right)\Bigg{]}$
@@
$\delta\,Z_{A^{\prime}}$ $=\frac{\alpha_{D}}{16\pi m_{A^{\prime}}^{4}(\epsilon-3)}\,\text{Re}\Bigg{[}m_{H}^{2}\left(m_{H}^{2}-2m_{A^{\prime}}^{2}\right)$ (30)
@@
$+m_{A^{\prime}}^{2}\Bigg{(}4m_{A^{\prime}}^{4}(\epsilon-3)+\left(m_{H}^{2}-m_{A^{\prime}}^{2}\right)A_{0}\left(m_{A^{\prime}}^{2}\right)$
$+4m_{A^{\prime}}^{2}m_{H}^{2}-m_{H}^{4}\Bigg{)}B_{0}^{\prime}\left(m_{A^{\prime}}^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)$
$+(m_{A^{\prime}}-m_{H})(m_{A^{\prime}}+m_{H})A_{0}\left(m_{H}^{2}\right)\Bigg{]}.$

### v2

@@ III.3 Relic density
Figure 4: Value of $\varepsilon$ producing the observed relic abundance of dark matter as a function of $m_{\chi}$, for $\alpha_{D}=$ 0.05, 0.3 and 0.5 (red, blue and magenta lines respectively) based on LO only (solid lines) and NLO + LO (dashed lines) cross sections and for the indicated value of $m_{A^{\prime}}/m_{\chi}$ and $m_{H}/m_{A^{\prime}}$ on each panel. For the $m_{A^{\prime}}=3m_{\chi}$ panel, the gray shaded regions are the exclusions from past accelerator experiments, and the green, purple and orange dotted lines are projections from the upcoming LDMX, $M_{3}$ and Belle II experiments respectively Gori et al. (2022). We show the experimental constraints in the $m_{A^{\prime}}=3m_{\chi}$ panel, which have been extensively studied in the literature. It is worth emphasizing that there are expected to be analogous constraints, on the other panels, but they require further analysis to be determined and will be addressed in future work.
In Fig. 4 we show plots of kinetic mixing as a function of DM mass for a few representative values of the ratio of $m_{A^{\prime}}/m_{\chi}$ and choose a benchmark parameter point $m_{H}=0.1m_{A^{\prime}}$. These plots were obtained by scanning through the kinetic mixing for each DM mass and finding the combination ($\varepsilon$, $m_{\chi}$) resulting in the observed relic abundance $\Omega h^{2}=0.12$ Aghanim et al. (2018). Each plot shows the required kinetic mixing parameter for three different values of the dark sector coupling $\alpha_{D}$, corresponding to the LO only (solid lines) and NLO + LO (dashed lines) computations, respectively. For $m_{A^{\prime}}/m_{\chi}=3$, which is typically chosen as a benchmark parameter point in the literature and is the most extensively studied. Hence, we show the current experimental constraints on this parameter space (a combination of both accelerator and astrophysical probes Gori et al. (2022)) in the gray shaded region. In that panel, the green, purple and orange dotted lines are sensitivity projections for the upcoming LDMX, $M_{3}$ and Belle II experiments, respectively. Assessing the experimental constraints on the other two panels is complex, and a thorough analysis is beyond the scope of this work. Following the trend in Fig. 3, we see in the top right panel, i.e. $m_{A^{\prime}}=2m_{\chi}$, the loop corrections result in not-so-significant changes in the thermal relic density lines, especially when lowering the value of the dark sector coupling constant $\alpha_{D}$. As one increases the value of $m_{A^{\prime}}/m_{\chi}$, we start seeing a significant change in the thermal relic line as a result of the loop corrections, especially for the large values of $\alpha_{D}$ which are often referenced in the literature, making these effects important to consider. We also note that the effect of these corrections can be even more dramatic if one further decreases the mass ratio $m_{H}/m_{A^{\prime}}$. Finally, in order for the future experiments to correctly characterize their future signals, the significance of these results should be taken into account.
@@ IV Dark matter direct detection
The spin averaged matrix element squared in the zero-momentum exchange limit is
$|\overline{\mathcal{M}_{EFT}}|^{2}=\Bigg{\{}\lvert\overline{\mathcal{M}}\rvert^{2}_{\text{LO}}+\overline{\delta\mathcal{M}^{2}}\Bigg{\}}_{t\rightarrow 0}$ (14)
where $\overline{\delta\mathcal{M}^{2}}$ corresponds to Eq. (9) and includes contributions from the one-loop corrections to the vertex, DM self-energy and the dark photon vacuum polarization, analogous to those represented in Fig. 2.
Given our focus on sub-GeV DM and the prospects for its detection at upcoming experiments, we restrict our attention to DM scattering with electrons. However our general computations can be extended to nucleon scattering with the incorporation of the appropriate form factors for either spin-dependent or spin-independent scattering.
The differential DM-electron scattering cross-section as a function of the momentum transfer $q$ is usually written
@@
$\overline{\sigma_{e}}=\overline{\sigma}_{e}^{\rm LO}+\overline{\delta\sigma_{e}}^{\rm NLO}_{\text{DM}}+\overline{\delta\sigma_{e}}^{\rm NLO}_{\text{Higgs}}+{\cal O}\left(\alpha_{D}^{3}\right).$ (16)
@@
$\overline{\delta\sigma_{e}}^{\text{NLO}}_{\text{DM}}$ $=\frac{\alpha_{D}\,\overline{\sigma_{e}}^{LO}}{12\,\pi\,m_{A^{\prime}}^{2}\,m_{\chi}^{2}}\Bigg{[}-m_{A^{\prime}}^{2}\left(-2\,m_{A^{\prime}}^{2}-4\,m_{\chi}^{2}\right)\Bigg{\{}4\,m_{\chi}^{2}B_{0}^{\prime}\left(m_{A^{\prime}}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)$ (18)
$-3\Bigg{(}B_{0}\left(0;m_{\chi}^{2},m_{\chi}^{2}\right)-B_{0}\left(m_{\chi}^{2};m_{A^{\prime}}^{2},m_{\chi}^{2}\right)$
$+\left(m_{A^{\prime}}^{2}-2\,m_{\chi}^{2}\right)C_{0}\left(0,m_{\chi}^{2},m_{\chi}^{2};m_{\chi}^{2},m_{\chi}^{2},m_{A^{\prime}}^{2}\right)-2\,m_{\chi}^{2}B_{0}^{\prime}\left(m_{\chi}^{2};m_{A^{\prime}}^{2},m_{\chi}^{2}\right)\Bigg{)}\Bigg{\}}$
$-16\,m_{\chi}^{4}\Bigg{(}B_{0}\left(m_{A^{\prime}}^{2};m_{\chi}^{2},m_{\chi}^{2}\right)-B_{0}\left(0;m_{\chi}^{2},m_{\chi}^{2}\right)\Bigg{)}\Bigg{]}$
The contribution to the scattering cross-section from the dark Higgs is
$\overline{\delta\sigma_{e}}^{\text{NLO}}_{\text{Higgs}}$ $=\frac{q_{\phi}^{2}\,\alpha_{D}\,\overline{\sigma_{e}}^{LO}}{6\,\pi\,m_{A^{\prime}}^{4}(m_{A^{\prime}}^{2}-m_{H}^{2})}\Bigg{[}(m_{A^{\prime}}^{2}-m_{H}^{2})\Bigg{\{}m_{A^{\prime}}^{2}\Bigg{(}\left(m_{A^{\prime}}^{2}-m_{H}^{2}\right)^{2}B_{0}^{\prime}\left(0;m_{A^{\prime}}^{2},m_{H}^{2}\right)$ (19)
$+\left(12\,m_{A^{\prime}}^{4}-4\,m_{A^{\prime}}^{2}\,m_{H}^{2}+m_{H}^{4}\right)B_{0}^{\prime}\left(m_{A^{\prime}}^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)\Bigg{)}$
$-2\left(6\,m_{A^{\prime}}^{4}-3\,m_{A^{\prime}}^{2}\,m_{H}^{2}+m_{H}^{4}\right)B_{0}\left(m_{A^{\prime}}^{2};m_{A^{\prime}}^{2},m_{H}^{2}\right)\Bigg{\}}$
$+2\left(6\,m_{A^{\prime}}^{4}-3\,m_{A^{\prime}}^{2}\,m_{H}^{2}+m_{H}^{4}\right)\Bigg{(}A_{0}\left(m_{A^{\prime}}^{2}\right)-\text{A}_{0}\left(m_{H}^{2}\right)\Bigg{)}\Bigg{]}$
where we have used the identity $B_{0}(0;m_{1}^{2},m_{2}^{2})=(A_{0}(m_{1}^{2})-A_{0}(m_{2}^{2}))/(m_{1}^{2}-m_{2}^{2})$.
@@
Figure 5: Dark matter direct detection parameter space. The left panel shows the ratio of NLO to LO electron scattering cross-section vs $m_{A^{\prime}}/m_{\chi}$ ratio for $m_{\chi}=1$ MeV and for $m_{H}/m_{A^{\prime}}=0.1,0.25,1\text{ and }10$, represented by the blue, magenta, red and orange lines, respectively. The plot for $m_{\chi}=1$ GeV is identical to the left panel. The right panel shows dark matter-electron scattering cross-section vs $m_{\chi}$ for dark matter thermal relic abundance with $\alpha_{D}=$, 0.05, 0.3 and 0.5, represented by the red, blue and magenta lines respectively, for $m_{H}/m_{A^{\prime}}=0.1$. We show both the LO only (solid lines) and NLO + LO (dashed lines) for the $m_{A^{\prime}}/m_{\chi}=2$ case. The upper gray shaded region represents the current model independent direct detection constraints from a combination of the XENON1T, PANDAX and SENSEI experiments and the vertical gray shaded region is the constraint from $\Delta N_{\text{eff}}$. All constraints were obtained from Ref. Essig et al. (2022).
@@
In Fig. 5, we show the direct detection constraints on the model. The left panel characterizes the importance of the one-loop correction, parametrized by the quantity $\overline{\delta\sigma}_{eNLO}/(\alpha_{D}\overline{\sigma}_{eLO})$ in terms of the ratio $m_{A^{\prime}}/m_{\chi}$ and for different values of $m_{H}/m_{A^{\prime}}$, similar to Fig. 3. Again, we find that this quantity is independent of the DM mass, and we choose $m_{\chi}=1$ MeV for reference.
@@
In the right panel, we show the averaged DM-electron scattering cross-section vs DM mass. The red, blue and magenta lines represent DM produced through the thermal freeze-out mechanism for benchmark choices of $\alpha_{D}=$ 0.05, 0.3 and 0.5 respectively. The dashed lines are NLO + LO contributions, while the solid lines are only at LO. For illustration, we only show the case for $m_{A^{\prime}}/m_{\chi}\simeq$ 2 and $m_{H}=0.1m_{A^{\prime}}$. We see a clear dependence on the dark sector coupling $\alpha_{D}$ and differences between the NLO + LO and LO only contributions.
@@
Also shown as the gray shaded region are the model independent constraints on this parameter space. Below $\sim$ 10 MeV, there is a stringent bound from big bang nucleosynthesis and $\Delta N_{\text{eff}}$, whereas the upper gray shaded region indicates bounds from a combination of direct detection experiments, including XENON1T, PANDAX and SENSEI (obtained from Ref. Essig et al. (2022)).
Similarly to the annihilation case above, we see the significance of the radiative corrections for different values of the ratio $m_{H}/m_{A^{\prime}}$, particularly for large $m_{A^{\prime}}/m_{\chi}$. For $m_{H}/m_{A^{\prime}}\geq 1$ the effect of a heavy dark Higgs is decreased. As one decreases $m_{H}/m_{A^{\prime}}$, we notice a significant correction to the total scattering cross-section that increases as $m_{A^{\prime}}/m_{\chi}$ rises, saturating for $m_{A^{\prime}}/m_{\chi}\sim 5$.
@@ V Discussion and Conclusions
In this work, we revisit the computations for sub-GeV thermal dark matter processes. We consider a highly motivated and sought after scenario, in which (pseudo-)Dirac dark matter couples to the SM via a massive dark photon, representing a compelling benchmark model of dark matter. We focus on the corrections of order $\alpha_{D}$, which is typically invoked to be large in order to obtain the observed relic density of the dark matter.
We compute the NLO corrections for thermal (pseudo-)Dirac dark matter annihilating in the early Universe, as this represents a very important target for upcoming experiments. We find that the NLO corrections can be quite significant, much larger than 10s of percents, depending on the $m_{A^{\prime}}/m_{\chi}$ ratio, the mass of the dark Higgs and the strength of the dark sector coupling constant $\alpha_{D}$. We find that this can result in a dramatic change compared to the targets inferred from tree-level calculations, as the mass of the Higgs in this model can be smaller compared to the mass of the dark photon. Hence it is an important result to establish the ability of proposed experiments to discover or constrain MeV scale dark matter, and to precisely quantify what we will learn from such experiments in the future. Furthermore, our results show that factorization into the “yield parameter” $Y$, as is common practice in the literature, can be misleading, especially in the resonance region, where the NLO corrections are more significant.
For completeness, we also show the direct detection limits based on DM-electron scattering. We find that the NLO corrections are as important as for the case of annihilation. One may wonder about the size of these corrections if the dark photon mass was generated through a different mechanism such as a Stueckelberg mass, and there were no dark Higgs particles. The scenarios for $m_{H}/m_{A^{\prime}}\geq 1$ in the plots above correspond to this case. As we see, the corrections are still noticeable, though not as significant as those obtained for lighter Higgs masses. We leave a more dedicated study of pseudo-Dirac dark matter with larger mass splittings as well as other Lorentz structures of interest to future studies.
@@ Acknowledgements
We thank Basudeb Dasgupta for useful discussions. G.M. acknowledges support from the UC office of the President through the UCI Chancellor’s Advanced Postdoctoral Fellowship. G.M. is also supported in part by the National Sciences and Engineering Research Council of Canada. T.M.P.T. and G.M. are supported in part by the U.S. National Science Foundation under Grant No. PHY-2210283. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation Grant No. PHY-2210452.
@@ B.2 Dark Higgs Contribution
$\Pi_{1}(k^{2})$ $=\frac{\alpha_{D}\,q_{\phi}^{2}}{4\,\pi\,(\epsilon-3)\,k^{2}}\Bigg{\{}\Bigg{(}-2k^{2}\left(m_{A}^{2}(2\epsilon-5)+m_{H}^{2}\right)$ (30)
$+k^{4}+\left(m_{A}^{2}-m_{H}^{2}\right)^{2}\Bigg{)}B_{0}\left(k^{2},m_{A}^{2},m_{H}^{2}\right)$
$+A_{0}\left(m_{H}^{2}\right)\left((5-2\epsilon)k^{2}+m_{A}^{2}-m_{H}^{2}\right)$
$-A_{0}\left(m_{A}^{2}\right)\left(k^{2}+m_{A}^{2}-m_{H}^{2}\right)\Bigg{\}}$
@@
$\delta m_{A^{\prime}}^{2}$ $=\frac{\alpha_{D}\,q_{\phi}^{2}}{4\,\pi\,m_{A^{\prime}}^{2}(\epsilon-3)}\text{Re}\Bigg{\{}\Bigg{(}-4m_{A^{\prime}}^{4}(\epsilon-3)$ (31)
@@
$-A\left(m_{H}^{2}\right)\left(2m_{A^{\prime}}^{2}(\epsilon-3)+m_{H}^{2}\right)$
$+\left(m_{H}^{2}-2m_{A^{\prime}}^{2}\right)A_{0}\left(m_{A^{\prime}}^{2}\right)\Bigg{\}}$
@@
$\delta\,Z_{A^{\prime}}$ $=\frac{\alpha_{D}\,q_{\phi}^{2}}{4\,\pi\,m_{A^{\prime}}^{4}(\epsilon-3)}\text{Re}\Bigg{\{}m_{H}^{2}\left(m_{H}^{2}-2m_{A^{\prime}}^{2}\right)$ (32)
@@
$+m_{A^{\prime}}^{2}\left(4m_{A^{\prime}}^{4}(\epsilon-3)+4m_{A^{\prime}}^{2}m_{H}^{2}-m_{H}^{4}\right)B_{0}^{\prime}\left(m_{A^{\prime}}^{2};\,m_{A^{\prime}}^{2},m_{H}^{2}\right)$
$+\left(m_{H}^{2}-m_{A^{\prime}}^{2}\right)A_{0}\left(m_{A^{\prime}}^{2}\right)$
$+(m_{A^{\prime}}-m_{H})(m_{A^{\prime}}+m_{H})A_{0}\left(m_{H}^{2}\right)\Bigg{\}}$